# Can Large Language Models be Effective Online Opinion Miners?

### v1

@@ 6 Conclusion
Despite its contributions, this study has several limitations, each of which also suggests promising directions for future research and practical extensions. First, although OOMB includes a diverse range of user-generated online content, it is currently confined to the vehicle domain, which may limit its generalizability to other areas such as electronics or healthcare. However, since the benchmark construction pipeline—including data collection, tuple annotation, and summary generation—is designed to be domain-agnostic, it can be easily extended to other fields with only minor adjustments to data sourcing and annotation guidelines.
@@
This study strictly adhered to ethical guidelines throughout the process of data collection and usage. Data crawling was conducted solely for non-commercial research purposes and performed at a controlled rate to avoid overloading servers or causing potential DDoS attacks. When collecting user reviews, personal information such as reviewer IDs, names, and locations was intentionally excluded, focusing only on text and dates to ensure user privacy. However, we cannot entirely rule out the possibility that the review text may contain personal details, hate speech, or inappropriate content. All data samples were collected and annotated in compliance with the terms and conditions of their respective sources. By making our dataset and models accessible, we aim to foster academic progress in generative event extraction research.
@@ A.1 Data Source
A.2 Human-in-the-loop Annotation Details
To ensure high-quality, consistent annotations, we adopted a human-in-the-loop process in which GPT-4o-mini 131313gpt-4o-mini-2024-07-18 serves as the initial annotator and human annotators [^14] We recruit undergraduates and graduates who are proficient in English and knowledgeable in the automotive domain. subsequently verify and refine its outputs. Prompts used to solicit these initial annotations are provided in Table 24 for tuples and Table 26 for summaries. All annotators underwent one full week of training on our detailed guidelines (Table 25 for tuples; Table 27 for summaries) before beginning any annotation work. This entire annotation–refinement process was applied to every sample in the dataset, ensuring higher overall quality and consistency in the resulting annotations.
Tuple verification and refinement were performed via our annotation UI (see Figure 8 and 9), which displays the original content, each tuple’s components with existence flags, and highlighted evidence sentences. Summary verification and refinement were conducted using a separate annotation UI (see Figure 10), which presents the content text, the associated gold tuple, and the working draft of the summary side by side for comparison and iterative improvement.
@@ A.4 Dataset Analysis
The t-SNE visualization shows that feature keywords form well-separated clusters according to major product aspects. Categories like Interior & Design, Driving Experience, and Performance & Powertrain appear frequently and show high cohesion. Engine & Driving Performance and Infotainment \$ Digital Systems are located near the core clusters, reflecting semantic proximity, while Overall Vehicle appears more scattered, indicating higher lexical variability. Outlier points on the periphery suggest rare or ambiguous feature expressions that may require finer handling.
@@ B.2 Feature-centric opinion extraction
As shown in Table 7 and 8, lower thresholds tend to excessively acknowledge partial overlaps, inflating recall to an unrealistic degree. Conversely, higher thresholds often miss semantically valid matches due to minor textual variations, causing the RM metric performance to converge toward Exact Match (EM) scores and consequently lose its intended flexibility. Thus, our experiments confirm that a threshold of 0.7 achieves optimal RM performance, which we subsequently adopt for our main experiments.
@@ B.3 Opinion-centric insight generation
LLM-Judge Evaluation Criteria
To evaluate summary quality across diverse criteria, we use GPT-4o and randomly sample 100 pieces of content from each of the four content types. The scoring scale for each evaluation follows the previous NLG evaluation framework, G-EVAL Liu et al. (2023a), and is measured on a scale from 1 to 5. We adopt the following six criteria:
  • Faithfulness: Evaluate whether the summary faithfully reflects the original review without distortion and check for any hallucinations.
  • Coverage: Evaluate whether the summary effectively captures and represents the key opinions expressed in the review.
  • Specificity: Evaluate whether the summary presents meaningful and relevant details rather than being vague or overly generic.
  • Insightfulness : Evaluate whether the summary provides meaningful insights that enhance understanding or decision-making for the reader.
  • Intent: Evaluate whether the summary accurately preserves the author’s original tone, intent, and nuances without altering the emotional or stylistic essence of the review.
  • Fluency: Evaluate whether the summary is naturally written, grammatically correct, and easy to read.
Case study: annotated opinion-centric summary
Table 28 presents representative examples of gold opinion-centric summaries for each content type.
Case study: LLMs’ summarize capability
We illustrate our LLM-Judge evaluation protocol with two case studies: one on a review site (Tables 29 and 31), and one on a Reddit (Tables 30 and 32). In each set, the first table shows the input content, the model-generated summary (GPT-4o), the gold opinion-centric summary, and the automatic metric A3CU score, while the second table provides a reference-free, six-dimension LLM-Judge assessment complete with per-dimension scores and detailed reasoning.
Although the A3CU metric assigns the review site example a high score (60.91) and the Reddit example a low score (10.12), our reference-free LLM-Judge evaluation reveals that the model’s performance on the Reddit content is in fact stronger across several human-aligned dimensions—particularly Coverage and Specificity. This divergence highlights the limitations of purely reference-based, automatic metrics in capturing the nuanced, insight-oriented qualities of opinion summaries. We therefore conclude that for OIG evaluation, combining automatic reference-based metrics with a reference-free, human-aligned judging protocol yields a more comprehensive and reliable assessment of LLMs’ true insight-generation capabilities.
@@
We assess the quality of the generated summary through a human evaluation conducted on Amazon Mechanical Turk (AMT). Specifically, we randomly sample 200 examples from our benchmark and ask three human judges per example to evaluate summaries generated by GPT-4o under two settings: 1) using only the input content, and 2) using both the input content and gold tuples. Each judge rates the quality of the summaries on a 1 to 5 scale across six criteria. The AMT interface used for human evaluation is presented in Figure 11, 12, 13, 14, 15, 16 and 17.
@@ B.4 Performance of LLMs by content type
We report the Feature-centric opinion extraction performance of various LLMs for each content type in Table 9, 10, 11 and 12 , and the Opinion-centric insight generation performance in Table 13, 14.
@@ B.5 Prompts
  • LLM-Judge Evaluation: The prompt for evaluating the LLM-Judge is presented in Table 18, 19, 20, 21, 22 and 23.
@@
You are a product and marketing manager at a global automotive company. Your task is to produce a free-form summary that categorizes and organizes a user text into higher-level insights, such that the report alone provides a clear understanding of the key opinions expressed. This summary should be written in natural, human-like language and structured around the core topics (features).
@@
4. Write the Summary Make it short (three to five lines). Use clear, direct language. Ensure that reading only this summary sufficiently conveys a user’s main viewpoints.
What to Avoid: Provide only the summary. Do not mention your process or how you arrived at it. Do not include introductory phrases such as “Here is a summary of the review” or “Based on the review text.” Do not directly copy sentences from the online text; rephrase and synthesize information. Carefully analyze the given text to determine the number of users and decide whether to use "user" or "users" accordingly. Note: Return your results in JSON format only, with the following structure: {’summary’: <str>}
@@
LLM-Judge Evaluation prompt - (Faithfulness)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the summary based on its Faithfulness. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Faithfulness (1-5) – Evaluate whether the summary accurately represents the original review without distorting its meaning, omitting key details, or introducing hallucinated information that was not present in the original text.
• Score 1: The summary completely distorts the original online content
@@
• Score 2: The summary significantly misrepresents the original online content with several incorrect information.
• Score 3: The summary partially reflects the original online content but has some minor incorrect information.
• Score 4: The summary significantly misrepresents the original online content with several incorrect information.
• Score 5: The summary completely reflects the original online content without any distortions and incorrect information.
@@
3. Evaluate whether the summary faithfully reflects the original review without any distortion.
4. Check if the summary contains any hallucinated information not present in the review.
@@
Table 18: The prompt for LLM-Judge Evaluation (Faithfulness).
LLM-Judge Evaluation prompt - (Coverage)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the summary based on its Coverage. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Coverage (1-5) – Evaluate whether the summary effectively captures and represents the key opinions expressed in the review without omitting significant details or key points. Ensure that all essential opinions are included and accurately represented.
•Score 1: The summary fails to capture any key opinions from the online content. The content is either completely missing or irrelevant to the original opinions.
•Score 2: The summary captures only a small portion of key opinions. Many important opinions from the online content are missing.
•Score 3: The summary captures some key opinions but misses others. The coverage is partial and could be more comprehensive.
•Score 4: The summary effectively captures most key opinions from the online content. The coverage is good but may miss minor details.
•Score 5: The summary comprehensively captures all key opinions from the online content. Nothing important is missing, and the coverage is complete.
@@
3. Evaluate whether the summary effectively captures and represents the key opinions expressed in the review.
4. Consider if any significant details or opinions are missing from the summary.
@@
Table 19: The prompt for LLM-Judge Evaluation (Coverage).
LLM-Judge Evaluation prompt - (Specificity)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the summary based on its Specificity. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Specificity (1-5) – Evaluate whether the summary presents meaningful and relevant details by including concrete information rather than being vague or overly generic. Ensure that the summary provides clear, detailed points that directly relate to the review content.
•Score 1: The summary is extremely vague and generic, lacking any meaningful details. It uses broad generalizations without specific examples or descriptions.
•Score 2: The summary includes very few specific details. Most information is presented in a general way without concrete examples.
•Score 3: The summary includes some specific details but could be more precise. There is a mix of specific and generic information.
•Score 4: The summary provides good specific details in most areas. The information is concrete and meaningful, though some minor points could be more detailed.
•Score 5: The summary is highly specific throughout, providing precise and meaningful details. All information is concrete with relevant examples and descriptions.
@@
3. Evaluate whether the summary provides specific, concrete details and avoids overly general or ambiguous statements.
@@
Table 20: The prompt for LLM-Judge Evaluation (Specificity).
LLM-Judge Evaluation prompt - (Insightfulness)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the summary based on its insightfulness. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Insightfulness (1-5) – Evaluate whether the summary provides meaningful insights that enhance understanding or decision-making for the reader. Ensure that the summary goes beyond a simple restatement of the review to offer unique interpretations or conclusions that add value.
•Score 1: The summary provides no meaningful insights. It simply restates basic facts without adding any value for understanding or decision-making.
•Score 2: The summary offers very limited insights. Most information is superficial and does not help readers gain deeper understanding.
•Score 3: The summary provides some useful insights but could go deeper. It offers moderate value for understanding and decision-making.
•Score 4: The summary provides good insights in most areas. The information is valuable for understanding and decision-making, though some points could be more insightful.
•Score 5: The summary provides excellent insights throughout. All information meaningfully enhances understanding and is highly valuable for decision-making.
@@
3. Evaluate whether the summary offers insightful, value-added interpretations that deepen understanding or guide decisions.
@@
Table 21: The prompt for LLM-Judge Evaluation (Insightfulness).
LLM-Judge Evaluation prompt - (Intent)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the summary based on its Intent. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Intent (1-5) – Evaluate whether the summary accurately preserves the author’s original tone, intent, and nuances without altering the emotional or stylistic essence of the review. Consider if the summary maintains the original sentiment and communication style throughout.
•Score 1: The summary completely fails to preserve the original tone and intent. The emotional essence and nuances are lost or significantly distorted.
•Score 2: The summary largely misrepresents the original tone and intent. Many nuances are missed or altered, though some basic sentiments remain intact.
•Score 3: The summary somewhat preserves the original tone and intent. Some nuances are captured while others are missed or altered.
•Score 4: The summary generally preserves the original tone and intent well. Most nuances and emotional elements are accurately captured, with only minor alterations.
•Score 5: The summary perfectly preserves the original tone, intent, and nuances. The emotional and stylistic essence is captured with complete accuracy.
@@
3. Evaluate whether the summary maintains the original tone, intent, and nuanced details of the review without altering its emotional or stylistic qualities.
@@
Table 22: The prompt for LLM-Judge Evaluation (Intent).
LLM-Judge Evaluation prompt - (Fluency)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the summary based on its Fluency. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Fluency (1-5) – Evaluate whether the summary is naturally written, grammatically correct, and easy to read. Consider whether the summary flows smoothly, uses proper grammar, and avoids awkward phrasing, ensuring it is accessible to the reader.
•Score 1: The summary is extremely difficult to read. It is filled with errors, awkward phrasing, and lacks proper grammar.
•Score 2: The summary is difficult to read. It contains many errors, awkward phrasing, and lacks proper grammar.
•Score 3: The summary is somewhat difficult to read. It has some errors, awkward phrasing, and lacks proper grammar.
•Score 4: The summary is generally easy to read. It has few errors, awkward phrasing, and lacks proper grammar.
•Score 5: The summary is extremely easy to read. It is filled with errors, awkward phrasing, and lacks proper grammar.
@@
3. Evaluate whether the summary is written in a natural and grammatically correct manner with clear and smooth flow.
@@
Table 23: The prompt for LLM-Judge Evaluation (Fluency).
@@
Opinion-centric-summary Annotation Prmopt
@@
You are a product and marketing manager at a global automotive company. You will be given a piece of user-generated automotive content, along with its final gold-standard set of (entity, feature, opinion) tuples. Your goal is to write a free-form opinion-centric summary that organizes and synthesizes the content into high-level, insightful categories.
Here is definition of (entity, feature, opinion) tuple and example: Entity: the brand and model of a vehicle for which an opinion is expressed (e.g., "volvo xc90", "toyota camry", "nissan sentra"). Feature: Feature: a specific characteristic, attribute, or component of the entity that users mention or evaluate. Includes aspects of performance, design, driving experience, features, systems, or parts (e.g., "ride comfort", "handling", "infotainment system", "interior", "brakes"). Opinion: Opinion: a subjective or objective evaluation, reaction, or judgment about a feature. The opinion span must be taken exactly from the text and contain no more than 5 words. If the feature is implicit (not explicitly stated in the text), label it as "NULL". This summary should be written in natural language, structured around the core features discussed, and clearly convey the most salient and recurring opinions.
Step 1. Read and Understand Thoroughly examine both the input content and the associated tuples. Identify the main opinion clusters and the intensity or frequency of key topics. Step 2. Organize into High-Level Topics Group similar tuples into broad categories (e.g., "Interior Design", "Performance & Handling"). Reflect on user sentiment and frequency to prioritize key points. Step 3. Generate a Three-to-Five-Line Summary Line 1: Highlight the most frequently mentioned or emphasized features. Line 2: Emphasize features that were praised or criticized with notable intensity or consensus. Line 3+: Provide a concluding sentence summarizing the overall sentiment or takeaway. Important Instructions The summary should be concise, natural, and informative—suitable for a marketing manager’s report Ensure it covers the major opinions expressed in the tuple set The tone should be neutral and professional, without exaggeration Do not directly copy sentences from the content Do not mention the annotation process or describe how the summary was generated Do not include phrases like “Based on the review…” or “Here is a summary…” Carefully analyze whether the input reflects a single user or multiple users, and adjust nouns/pronouns accordingly ("user" vs. "users"). Note: Return your results in JSON format only, with the following structure: {’summary’: <str>}
@@
Table 26: OOMB: Opinion-centric summary annotation prompt.
Opinion-centric-summary Annotation & Verification Guideline

### v2

@@ 6 Conclusion
Despite its contributions, this study has several limitations, each of which also suggests promising directions for future research and practical extensions. First, although OOMB includes a diverse range of user-generated online content, it is currently confined to the vehicle domain, which may limit its generalizability to other areas such as electronics or healthcare. However, since the benchmark construction pipeline, which includes data collection, tuple annotation, and insight generation, is designed to be domain-agnostic, it can be easily extended to other fields with only minor adjustments to data sourcing and annotation guidelines.
@@
This study strictly adhered to ethical guidelines throughout the entire process of data collection and usage. As the benchmark is limited to the automotive domain, concerns may arise regarding representativeness and potential amplification of biases related to gender, socioeconomic status, or geography. To address this issue, we collected data from global platforms such as Reddit, YouTube and well-known review blogs which are widely used by diverse communities across different regions and demographics. We believe this design helps capture a broad spectrum of opinions present in real-world online discourse, though we acknowledge that residual biases may still persist.
Regarding data collection procedures, crawling was conducted solely for non-commercial research purposes and performed at a controlled rate to prevent server overload or potential DDoS attacks. When collecting user reviews, personal information such as reviewer IDs, names, and locations was intentionally excluded, focusing only on textual content to ensure user privacy. However, we cannot completely rule out the possibility that review texts may contain sensitive personal details, hate speech, or inappropriate content.
Finally, all data samples were collected and annotated in compliance with the terms and conditions of their respective sources. By releasing our dataset, we aim to contribute to the academic advancement of generative opinion mining research.
Acknowledgments
This work was supported by the IITP grants funded by the Korea government (MSIT) (No. RS-2020- II201361; RS-2024-00457882, AI Research Hub Project), and the NRF grant funded by the Korea government (MSIT) (No. RS-2025-00560295).
@@ A.1 Data Source
A.2 Data Annotation Details
Human-in-the-loop annotation
To ensure high-quality, consistent annotations, we adopted a human-in-the-loop process in which GPT-4o-mini 151515gpt-4o-mini-2024-07-18 serves as the initial annotator and human annotators [^16] We recruit undergraduates and graduates who are proficient in English and knowledgeable in the automotive domain. subsequently verify and refine its outputs. Prompts used to solicit these initial annotations are provided in Table 25 for tuples and Table 27 for insights. All annotators underwent one full week of training on our detailed guidelines (Table 26 for tuples; Table 28 for insights) before beginning any annotation work. This entire annotation–refinement process was applied to every sample in the dataset, ensuring higher overall quality and consistency in the resulting annotations.
Inter-Annotator Agreement Details
We provide additional details of the inter-annotator agreement evaluation reported in Table 2.
  • Span-level tuple agreement: Each annotator independently extracted (entity, feature, opinion) tuples from every sentence in the selected contents. Inter-annotator consistency was measured using Fleiss’ Kappa, yielding a score of 0.6821, which generally indicates substantial agreement among multiple annotators. This result indicates strong inter-annotator reliability in capturing core opinion structures at the span level.
  • Preferred insight agreement: For each content, annotators selected their top-1 to top-3 preferred opinion-centric insights from five LLM-generated candidate insights. We computed Fleiss’ Kappa over these preference rankings and obtained a score of 0.7695, indicating substantial consistency among annotators.
  • Insight semantic consistency: After guideline-based verification and rewriting, we computed the pairwise semantic similarity between the finalized opinion-centric insights produced by different annotators. We used BERTScore to measure semantic alignment, and report the average score across all annotator pairs, which reached a very high value of 0.9297. This indicates that, despite differences in wording, annotators consistently expressed the same underlying meaning.
Annotation Tools
To enhance clarity and efficiency in the annotation process, we developed dedicated UIs for each stage of verification and refinement. Tuple verification and refinement were performed via our annotation UI (see Figure 8 and 9), which displays the original content, each tuple’s components with existence flags, and highlighted evidence sentences. Insight verification and refinement were conducted using a separate annotation UI (see Figure 10), which presents the content text, the associated gold tuple, and the working draft of the insight side by side for comparison and iterative improvement.
@@ A.4 Dataset Analysis
The t-SNE visualization shows that feature keywords form well-separated clusters according to major product aspects. Categories like Interior & Design, Driving Experience, and Performance & Powertrain appear frequently and show high cohesion. Engine & Driving Performance and Infotainment & Digital Systems are located near the core clusters, reflecting semantic proximity, while Overall Vehicle appears more scattered, indicating higher lexical variability. Outlier points on the periphery suggest rare or ambiguous feature expressions that may require finer handling.
@@ B.2 Feature-centric opinion extraction
As shown in Table 8 and 9, lower thresholds tend to excessively acknowledge partial overlaps, inflating recall to an unrealistic degree. Conversely, higher thresholds often miss semantically valid matches due to minor textual variations, causing the RM metric performance to converge toward Exact Match (EM) scores and consequently lose its intended flexibility. Thus, our experiments confirm that a threshold of 0.7 achieves optimal RM performance, which we subsequently adopt for our main experiments.
@@ B.3 Opinion-centric insight generation
LLM-as-a-judge evaluation criteria
To evaluate insight quality across diverse criteria, we use GPT-4o and randomly sample 100 pieces of content from each of the four content types. The scoring scale for each evaluation follows the previous NLG evaluation framework, G-EVAL Liu et al. (2023a), and is measured on a scale from 1 to 5. We adopt the following six criteria:
  • Faithfulness: Evaluate whether the insight faithfully reflects the original review without distortion and check for any hallucinations.
  • Coverage: Evaluate whether the insight effectively captures and represents the key opinions expressed in the review.
  • Specificity: Evaluate whether the insight presents meaningful and relevant details rather than being vague or overly generic.
  • Insightfulness : Evaluate whether the insight offers meaningful observations or interpretations that enhance understanding or decision-making for the reader.
  • Intent: Evaluate whether the insight accurately preserves the author’s original tone, intent, and nuances without altering the emotional or stylistic essence of the review.
  • Fluency: Evaluate whether the insight is naturally written, grammatically correct, and easy to read.
Case study: Annotated opinion-centric insight
Table 29 presents representative examples of gold opinion-centric insights for each content type.
Case study: LLMs’ insight generation capability
We illustrate our LLM-as-a-judge evaluation protocol with two case studies: one on a review site (Tables 30 and 32), and one on a Reddit (Tables 31 and 33). In each set, the first table shows the input content, the model-generated insight (GPT-4o), the gold opinion-centric insight, and the automatic metric A3CU score, while the second table provides a reference-free, six-dimension LLM-as-a-judge assessment complete with per-dimension scores and detailed reasoning.
Although the A3CU metric assigns the review site example a high score (60.91) and the Reddit example a low score (10.12), our reference-free LLM-as-a-judge evaluation reveals that the model’s performance on the Reddit content is in fact stronger across several human-aligned dimensions—particularly Coverage and Specificity. This divergence highlights the limitations of purely reference-based, automatic metrics in capturing the nuanced, insight-oriented qualities of opinion insights. We therefore conclude that for OIG evaluation, combining automatic reference-based metrics with a reference-free, human-aligned judging protocol yields a more comprehensive and reliable assessment of LLMs’ true insight-generation capabilities.
@@
We assess the quality of the generated insight through a human evaluation conducted on Amazon Mechanical Turk (AMT). Specifically, we randomly sample 200 examples from our benchmark and ask three human judges per example to evaluate insights generated by GPT-4o under two settings: 1) using only the input content, and 2) using both the input content and gold tuples. Each judge rates the quality of the insights on a 1 to 5 scale across six criteria. The AMT interface used for human evaluation is presented in Figure 11, 12, 13, 14, 15, 16 and 17.
@@ B.4 Performance of LLMs by content type
We report the Feature-centric opinion extraction performance of various LLMs for each content type in Table 10, 11, 12 and 13 , and the Opinion-centric insight generation performance in Table 14, 15.
@@ B.5 Prompts
  • LLM-as-a-judge Evaluation: The prompt for evaluating the LLM-Judge is presented in Table 19, 20, 21, 22, 23 and 24.
@@
You are a product and marketing manager at a global automotive company. Your task is to produce a free-form insight text that categorizes and organizes a user text into higher-level insights, such that the report alone provides a clear understanding of the key opinions expressed. This insight should be written in natural, human-like language and structured around the core topics (features).
@@
4. Write the Insight Make it short (three to five lines). Use clear, direct language. Ensure that reading only this insight sufficiently conveys a user’s main viewpoints.
What to Avoid: Provide only the insight text. Do not mention your process or how you arrived at it. Do not include introductory phrases such as “Here is a insight of the review” or “Based on the review text.” Do not directly copy sentences from the online text; rephrase and synthesize information. Carefully analyze the given text to determine the number of users and decide whether to use "user" or "users" accordingly. Note: Return your results in JSON format only, with the following structure: {’insight’: <str>}
@@
LLM-as-a-judge Evaluation prompt - (Faithfulness)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the insight based on its Faithfulness. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Faithfulness (1-5) – Evaluate whether the insight accurately represents the original review without distorting its meaning, omitting key details, or introducing hallucinated information that was not present in the original text.
• Score 1: The insight completely distorts the original online content
@@
• Score 2: The insight significantly misrepresents the original online content with several incorrect information.
• Score 3: The insight partially reflects the original online content but has some minor incorrect information.
• Score 4: The insight significantly misrepresents the original online content with several incorrect information.
• Score 5: The insight completely reflects the original online content without any distortions and incorrect information.
@@
3. Evaluate whether the insight faithfully reflects the original review without any distortion.
4. Check if the insight contains any hallucinated information not present in the review.
@@
Table 19: The prompt for LLM-as-a-judge Evaluation (Faithfulness).
LLM-as-a-judge Evaluation prompt - (Coverage)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the insight based on its Coverage. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Coverage (1-5) – Evaluate whether the insight effectively captures and represents the key opinions expressed in the review without omitting significant details or key points. Ensure that all essential opinions are included and accurately represented.
•Score 1: The insight fails to capture any key opinions from the online content. The content is either completely missing or irrelevant to the original opinions.
•Score 2: The insight captures only a small portion of key opinions. Many important opinions from the online content are missing.
•Score 3: The insight captures some key opinions but misses others. The coverage is partial and could be more comprehensive.
•Score 4: The insight effectively captures most key opinions from the online content. The coverage is good but may miss minor details.
•Score 5: The insight comprehensively captures all key opinions from the online content. Nothing important is missing, and the coverage is complete.
@@
3. Evaluate whether the insight effectively captures and represents the key opinions expressed in the review.
4. Consider if any significant details or opinions are missing from the insight.
@@
Table 20: The prompt for LLM-as-a-judge Evaluation (Coverage).
LLM-as-a-judge Evaluation prompt - (Specificity)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight insight that groups user opinions at the topic level to offer insights. Your task is to evaluate the insight based on its Specificity. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Specificity (1-5) – Evaluate whether the insight presents meaningful and relevant details by including concrete information rather than being vague or overly generic. Ensure that the insight provides clear, detailed points that directly relate to the review content.
•Score 1: The insight is extremely vague and generic, lacking any meaningful details. It uses broad generalizations without specific examples or descriptions.
•Score 2: The insight includes very few specific details. Most information is presented in a general way without concrete examples.
•Score 3: The insight includes some specific details but could be more precise. There is a mix of specific and generic information.
•Score 4: The insight provides good specific details in most areas. The information is concrete and meaningful, though some minor points could be more detailed.
•Score 5: The insight is highly specific throughout, providing precise and meaningful details. All information is concrete with relevant examples and descriptions.
@@
3. Evaluate whether the insight provides specific, concrete details and avoids overly general or ambiguous statements.
@@
Table 21: The prompt for LLM-as-a-judge Evaluation (Specificity).
LLM-as-a-judge Evaluation prompt - (Insightfulness)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the insight based on its insightfulness. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Insightfulness (1-5) – Evaluate whether the insight provides meaningful insights that enhance understanding or decision-making for the reader. Ensure that the insight goes beyond a simple restatement of the review to offer unique interpretations or conclusions that add value.
•Score 1: The insight provides no meaningful insights. It simply restates basic facts without adding any value for understanding or decision-making.
•Score 2: The insight offers very limited insights. Most information is superficial and does not help readers gain deeper understanding.
•Score 3: The insight provides some useful insights but could go deeper. It offers moderate value for understanding and decision-making.
•Score 4: The insight provides good insights in most areas. The information is valuable for understanding and decision-making, though some points could be more insightful.
•Score 5: The insight provides excellent insights throughout. All information meaningfully enhances understanding and is highly valuable for decision-making.
@@
3. Evaluate whether the insight offers insightful, value-added interpretations that deepen understanding or guide decisions.
@@
Table 22: The prompt for LLM-as-a-judge Evaluation (Insightfulness).
LLM-as-a-judge Evaluation prompt - (Intent)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the insight based on its Intent. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Intent (1-5) – Evaluate whether the insight accurately preserves the author’s original tone, intent, and nuances without altering the emotional or stylistic essence of the review. Consider if the insight maintains the original sentiment and communication style throughout.
•Score 1: The insight completely fails to preserve the original tone and intent. The emotional essence and nuances are lost or significantly distorted.
•Score 2: The insight largely misrepresents the original tone and intent. Many nuances are missed or altered, though some basic sentiments remain intact.
•Score 3: The insight somewhat preserves the original tone and intent. Some nuances are captured while others are missed or altered.
•Score 4: The insight generally preserves the original tone and intent well. Most nuances and emotional elements are accurately captured, with only minor alterations.
•Score 5: The insight perfectly preserves the original tone, intent, and nuances. The emotional and stylistic essence is captured with complete accuracy.
@@
3. Evaluate whether the insight maintains the original tone, intent, and nuanced details of the review without altering its emotional or stylistic qualities.
@@
Table 23: The prompt for LLM-as-a-judge Evaluation (Intent).
LLM-as-a-judge Evaluation prompt - (Fluency)
@@
You are provided with an online user’s automobile review text along with an opinion-centric insight summary that groups user opinions at the topic level to offer insights. Your task is to evaluate the insight based on its Fluency. Make sure you understand the following evaluation metric very clearly.
Evaluation Criteria: Fluency (1-5) – Evaluate whether the insight is naturally written, grammatically correct, and easy to read. Consider whether the insight flows smoothly, uses proper grammar, and avoids awkward phrasing, ensuring it is accessible to the reader.
•Score 1: The insight is extremely difficult to read. It is filled with errors, awkward phrasing, and lacks proper grammar.
•Score 2: The insight is difficult to read. It contains many errors, awkward phrasing, and lacks proper grammar.
•Score 3: The insight is somewhat difficult to read. It has some errors, awkward phrasing, and lacks proper grammar.
•Score 4: The insight is generally easy to read. It has few errors, awkward phrasing, and lacks proper grammar.
•Score 5: The insight is extremely easy to read. It is filled with errors, awkward phrasing, and lacks proper grammar.
@@
3. Evaluate whether the insight is written in a natural and grammatically correct manner with clear and smooth flow.
@@
Table 24: The prompt for LLM-as-a-judge Evaluation (Fluency).
@@
Opinion-centric-insight Annotation Prmopt
@@
You are a product and marketing manager at a global automotive company. You will be given a piece of user-generated automotive content, along with its final gold-standard set of (entity, feature, opinion) tuples. Your goal is to write a free-form opinion-centric insight summary that organizes and synthesizes the content into high-level, insightful categories.
Here is definition of (entity, feature, opinion) tuple and example: Entity: the brand and model of a vehicle for which an opinion is expressed (e.g., "volvo xc90", "toyota camry", "nissan sentra"). Feature: Feature: a specific characteristic, attribute, or component of the entity that users mention or evaluate. Includes aspects of performance, design, driving experience, features, systems, or parts (e.g., "ride comfort", "handling", "infotainment system", "interior", "brakes"). Opinion: Opinion: a subjective or objective evaluation, reaction, or judgment about a feature. The opinion span must be taken exactly from the text and contain no more than 5 words. If the feature is implicit (not explicitly stated in the text), label it as "NULL". This insight should be written in natural language, structured around the core features discussed, and clearly convey the most salient and recurring opinions.
Step 1. Read and Understand Thoroughly examine both the input content and the associated tuples. Identify the main opinion clusters and the intensity or frequency of key topics. Step 2. Organize into High-Level Topics Group similar tuples into broad categories (e.g., "Interior Design", "Performance & Handling"). Reflect on user sentiment and frequency to prioritize key points. Step 3. Generate a Three-to-Five-Line Insight Line 1: Highlight the most frequently mentioned or emphasized features. Line 2: Emphasize features that were praised or criticized with notable intensity or consensus. Line 3+: Provide a concluding sentence summarizing the overall sentiment or takeaway. Important Instructions The insight should be concise, natural, and informative—suitable for a marketing manager’s report Ensure it covers the major opinions expressed in the tuple set The tone should be neutral and professional, without exaggeration Do not directly copy sentences from the content Do not mention the annotation process or describe how the insight was generated Do not include phrases like “Based on the review…” or “Here is a insight…” Carefully analyze whether the input reflects a single user or multiple users, and adjust nouns/pronouns accordingly ("user" vs. "users"). Note: Return your results in JSON format only, with the following structure: {’insight’: <str>}
@@
Table 27: OOMB: Opinion-centric insight annotation prompt.
Opinion-centric-insight Annotation & Verification Guideline